OpenStudy (anonymous):

A container holds 7/10 of water . After pouring 6 gallons of water in it it is now 1/2 full . How many gallons can the container hold ? @kelliegirl33

OpenStudy (anonymous):

7/10 - 1/2 = 7/10 - 5/10 = 2/10 = 1/5 6 gallons is 1/5 so if 6 gallons is 1/5, then a full tank is 30 gallons

OpenStudy (anonymous):

Thank youu! can u help me with one more!

OpenStudy (anonymous):

I can try